PHOS commissioning in 2010

24 June 2010 PHOS Commissioning Meeting

Minutes

- Point 2

Most of the lost raw data from the 900 GeV run has been recovered by the CERN IT department. There has been no data taking with stable beams since 5. June because of machine commissioning. Physics data taking is expected to resume tonight. All subsystems have been requested to provide a Heavy-Ion like configuration for tests (i.e. a configuration which provides approximately the same event size as in heavy-ion collisions). PHOS has done this and has taken part in several technical runs with this configuration.

- RCU

The connection to Branch A, RCU0 on Module 3 was lost after the access during the technical stop 31/5 - 2/6. Presumably someone went under Module 3 and touched the RCU which led to it being disconnected. This was fixed during an access on 8/6 (Sasha and Oystein). The same problem appeared again yesterday, again following an access into L3 on Tuesday. It has still not been recovered. Access is needed.

- DCS

The HV server was down for 3 days without anyone noticing. It should be part of the daily routine of the PHOS On Call to check this connection at least once daily.

- Trigger

The TRU fakealtro data from the test setup in Bergen can be read back. There are problems with generating a signal in the test setup so the timing issue has not yet been studied.

- LTU

The LTU lent to the TPC people has been returned.

- PHOS Shifter/OnCall manual

A new version is available from the ALICE Run pages (Joakim, Sergey).

27 May 2010 PHOS Commissioning Meeting

Minutes

- Point 2

The raw data from the 900 GeV runs taken about 3 weeks ago have been lost. New 900 GeV runs are being taken now. There will be a technical stop next week 31/5 - 2/6.

- Shifts

The PHOS shifts were cancelled from 17/5. No problems encountered so far.

- Electronics

A new RCU firmware version from Attiq is available. It will be installed and tested during the technical stop next week. It should hopefully reduce the dead-time significantly.

- Trigger

A global run (120967) with PHOS triggered by PHOS when the LED system was running was taken yesterday. The fakealtro data were written out. An analysis of this data did not show any signal in the fakealtro data, however. This is being investigated by Lijiao and Dong.

- DCS

There was a problem with the LED VME crate during last weekend. What caused it is unknown. The problem fixed itself It is being followed up by Alexandre Vinogradov.

- LTU

No news.

6 May 2010 PHOS Commissioning Meeting

Minutes

- Point 2

Running at 900 GeV during the previous weekend. 7 M events were recorded. The plans for LHC are to do machine development during the day and physics running (7 TeV) during the night.

- DCS

Alexander Mamonov has been working on the implementation of the LED system in PVSS. He is also working on updating the text for the PHOS alarms and improving the PHOS alarms mailing list. At the moment, only an interlock for the cooling plant generates a mail, but in the future any hardware or software interlock will send a mail to the alarm list. The changes to the FEE control done by Kyrre last week will also be included in the PVSS.

- Shifts

Progress on writing of documentation for global shifters and PHOS On Call:

DCS (Alexander Mamonov) - in progress; DQM (Oystein, Yuri Kharlov) - in progress; On Call (Joakim) - Included in the new version of the shift manual.

Sergey Nikolaev is working on updating the PHOS manual to incorporate the latest changes to the DCS system.

- LTU

No news. Bernhard will check.

29 april 2010 PHOS Commissioning Meeting

Minutes

- Point 2

20 Mevents were recorded last weekend. The total number of recorded events is now more than 35 M. One run with 100,000 physics triggers has been taken with the PHOS TRUs being read out. There has been a technical stop Mon-Wed this week. The LHC is recovering and the first stable beams are expected Saturday evening. The goal is then to collect 10 Mevents at 900 GeV before continuing running at 7 TeV.

- DCS

Henrik and Kyrre have been working on implementation of FEE control this week. It is now possible to configure from PVSS, but the actual configuration commands are not yet implemented. The temperatures can only be read from about half of the FEE cards. Alexander Mamonov has also made several updates to the DCS systems but since he is not present at the meeting the details are not available.

- Shifts

We still keep PHOS shifts during day and evening. A set of instructions for global DCS shifter (Alexander Mamonov), DQM shifter (Yuri Kharlov, Oystein Djuvsland), PHOS On Call (Joakim Nystrand) have to be written before the shifts can be cancelled. The current plan is to cancel shifts from 15. May. Alexei Kuryakin has created a mailing list for notification in case of an cooling interlock.

- Cooling system

The cooling system has been reconnected to the UPS. The reset button in CR-4 has been checked.

15 april 2010 PHOS Commissioning Meeting

Minutes

- Point 2

Commissioning of the accelerator during the day; physics running during the night. PHOS remains stable in the runs.

- Trigger

Tests of trigger timing module by module. Two TRUs have had to be masked out.

- DCS

Alexander Mamonov is now at CERN until 16 May. He will work on improving the PHOS DCS and PHOS in the Alice Configuration Tool (ACT). Kyrre and Henrik will go to CERN for the technical stop 26-28 April to work on the Intercom layer.

- Shifts

The proposal to end PHOS shifts by 1. May is being considered.

8 april 2010 PHOS Commissioning Meeting

Minutes

- Point 2

First pp collisions at 7 TeV occured on 30 March. Since then about 8 Mevents have been accumulated. PHOS has been very stable and has participated in all runs taken. Today there is access. More collisions are expected on Saturday and Sunday.

- Gain settings

The first analysis of the new data at 7 TeV taken with fixed pedestals indicate that the current calibration is off by a factor 2.5. We have to decide as soon as possible if we should keep the current gain or reduce it.

- Trigger

The new trigger firmware was tested on Sunday. The signal from PHOS now arrives earlier than the other subsystems. There is a larger jitter now (3-4 beam clocks) than before. This is not expected to be a serious problem but Anton will be consulted.

See timing plots here

orbit1

orbit2

- DCS

There remains work to be done on ACT and on implementation of control of FEE in PVSS. The LHC schedule should be investigated so that people can plan their travel to CERN.

- Shifts

A discussion of when it is time to cancel shifts should be started. At the last PHOS meeting it was proposed to end shifts at the end of April. It is important that experts sign up for On Call shifts for PHOS.

- PHOS lab

The missing LTU is still missing.

18 march 2010 PHOS Commissioning Meeting

Minutes

- Point 2

Cosmic data have been accumulated with the TRU trigger during the last week. About 700,000 events have been recorder. Collisions at 900 GeV now expected on Sunday (21/3), starting at midnight 00:00.

- Pulse shape

Problem with 2-peak structure has been solved by introducing fixed pedestals. Luciano has an explanation why the two-peak structure can appear if the automatic pedestal calculation is interrupted by a signal.

The stability of the pedestal has been checked. The difference between pedestals taken 3 days apart are very small; the mean has changed by roughy 0.1 unit or less depending on Module and the RMS is also low, roughly 0.1 unit or less. There are some outliers, but those are all associated with channels that are already in the bad-channel map.

See Oystein's presentation here http://indico.cern.ch/getFile.py/access?resId=1&materialId=slides&confId=88546

- Results from cosmic runs

Results from Boris. Good matching with TPC tracks, = 2.5 cm. The energy spectrum shows a nice mip-like peak with a mean, however, at 770 MeV. This would correspond to a APD gain of about 150 rather than the expected 50. This is not understood.

See Boris' plots here http://indico.cern.ch/getFile.py/access?resId=0&materialId=slides&confId=88546

- Trigger

Because of the work with the pulse shape and implementation of fixed pedestals it has not been possible to test the new trigger firmware. It should be done as soon as possible.

- DCS

The ACT problem for LED runs persists.

- Shifts/OnCall

The night shifts have been replaced by OnCall. Most subsystems, most likely also PHOS, will cancel all shifts from mid-April. People are encouraged to get trained and sign up for shifts for the global systems like DCS.

- PHOS lab

An LTU is still missing. Oystein will look into it. A new VME crate is needed. Iouri will change it.

11 march 2010 PHOS Commissioning Meeting

Minutes

- Point 2

Beams are circulating in the machine most of the time. Collisions at 900 GeV are planned for 3 shifts on Friday 19 March. Collisions at 7 TeV are expected before the end of this month.

- Trigger

A new version has been developed which meets the 800 ns time limit for L0 input to the CTP. The time sum has been removed and the signal is extracted from a single 50ns time bin. Space sum is still performed.

- DCS

There is still a small memory leak present when opening the LV panels. There are still problems with handling LED runs in the ACT (ALICE Configuration Tool).

- Raw data

Dmitri reported on the details of the problem with corrupted data when PHOS is triggered by the CTP. His report is available here

http://indico.cern.ch/conferenceDisplay.py?confId=87896

Oystein described the progress made in implementing reading of fixed pedestals and loading them into the ALTRO. It is expected that this will fix the corruption problem.

- PHOS lab

Date has been updated on the pc in the PHOS lab. One LTU is still missing.

- Shifts

The need for people to serve as PHOS OnCall was reiterated.

4 march 2010 PHOS Commissioning Meeting

Minutes

- Point 2

Circulating beams are expected next week.

- Trigger

The TED test last Friday showed that the PHOS trigger works but that it is delayed by 5-6 beam clocks (about 150 ns). Firmware with a new trigger algorithm is being developed to make it faster.

It has been found that LED signals and signals from Cosmics taken with the PHOS trigger are corrupted and show a two-peak structure rather than the expected Gamma-2 shape. The cause for this is being investigated.

- DCS

Thanks to the work of Alexander, Kyrre, and Jiro several problems have been fixed. The memory leak when opening the LV panel is now gone. The fluctuating HV states have been fixed by changing the FSM logic.

- FEE

The PC aldaqpc019 in the PHOS lab will be upgraded and a new version of date (DAQ) will be installed.

- Shifts

Since PHOS runs stably and there is not much to do during the night, all or most night shifts can be cancelled and replaced by having people OnCall.

- Cooling system

The cooling system will be removed from the UPS; only the control system will be connected to UPS.

25 Feb 2010 PHOS Commissioning Meeting

Minutes

- Point 2

Last Thursday evening there were beam splashes on the TED. PHOS took data. More splashes are expected tomorrow, Friday 26/2.

- Trigger

The tests of the PHOS trigger during the TED dumps were inconclusive because of secondary triggers generated by the read out. This was fixed on Sunday. Since then, cosmics have been collected with the PHOS trigger during the nights. On Wednesday (24/2) PHOS was for the first time used to trigger the ALICE experiment in a clusted which included PHOS and the TPC.

- PVSS

Kyrre will work with Marco on Monday to solve the remaining ACT problems. The problems with fluctuating (HV) states have been traced to the ISEGs. One feature is that the clock is shifted by 15 min per day. A memory leak has been spotted which leads to a frozen LV display if it is kept open for more than 1-2 mins.

- FEE

Repair of FEEs in PHOS lab. The PC in the phos lab (aldaqpc019) has to be upgraded to SLC5 in order to get the latest version of DAQ (date). Users who want to save their data on the disk should make backups.

18 Feb 2010 PHOS Commissioning Meeting

Minutes

- Trigger

6 good TRUs yesterday, today 1 or 2.

m4 TRUs un-configured state after data taking

LED only one module?

- Cooling

stable

- RCU

- IntercomLayer

New java icl is installed, works better than the old one. Need to reduce the number of services. Re-programming the RCU not working with new FeeServer.

"no link" seen in PVSS.

- PVSS

ALICE Configuration Tool test. LED - not working, PHYSICS - working. Kyrre working on solutions to problems.

11 Feb 2010 PHOS Commissioning Meeting

Minutes

- Point 2

The L3 magnet will close at 17:00 today and the field will be turned on at 18:00. The ALICE cavern will close on 15. February. PHOS would like to participate if there are runs with dumps on TED. This will probably happen next week.

- Trigger

Lijiao presented results from the trigger tests. When the LED system is run at 20 Hz one can read back a stable trigger rate of 20 Hz from TOR. The background rate is below 1 Hz. The trigger threshold is set at x100. The influence of a possible LED phase shift between modules was discussed but no firm conclusion was reached. The next step in the trigger commissioning will be to set up a global partition with a PHOS trigger from the CTP. Lijiao's presentation is here http://indico.cern.ch/materialDisplay.py?materialId=slides&confId=84724

- LED

Most of module 3 has now been illuminated by the LED system but there are still some bugs in the LED firmware or software.

- Cooling

Monday, Wednesday, Thursday this week there have been interventions to the PHOS crystal cooling plant. On Monday a broken power supply was replaced. Hopefully the intervention today will be the last one before the closing of the ALICE cavern. In the future, these interventions have to be announced further in advance, for example at the commissioning meetings.

- RCU

The PHOS busy time is now 830 microsec, down from 1400 microsec last year. There is hope that it can be reduced by another 100-200 microsecs through upgrades of the RCU firmware.

- FEE

Bernhard and Iouri will start repairing the 28 broken spare FEEs on Monday.

- AMORE

The functionality should be tested.

- Intercom layer

A new java-based version of the Intercom layer has been installed by Kyrre with help from Oystein and Timur. It remains to get the database lists in correctly and it still kills the FEE servers occasionally,

- PVSS

The PHOS PVSS system stores too much information on temperatures and voltages. In fact, the volume from PHOS is larger than from all other subsystems combined. This must be reduced. The best way would be to implement smoothing, i.e. only save a value if it deviates significantly from the previous reading. Kyrre, Serguei, Yuri, and Alexander are looking into this.

28 Jan 2010 PHOS Commissioning Meeting

Minutes

- Point 2

The main activity during the past week has been trigger develpoment work done from Bergen. From Wednesday afternoon around 4 PM PHOS has been off because of the UPS intervention.

- Trigger/TRU

The tests by Jiri, Dong, and Lijiao on the test bench in Bergen have been successful. The next step is to take pedestal data at Point 2 and find a map of noisy channels to mask out. Then one can test to trigger on a LED signal.

- LED

LED data have been taken for most parts of module 3. The missing parts are due to a bug in the LED software which should be fixable. From LED data from Module 2, Yuri has found that the some of the APDs in that module appear to be overcalibrated.

- Sparse read-out

Sparse read-out now works for EmCal, but improvements in the RCU firmware for increased speed is still being worked on. The implementation of sparse read-out in EmCal only involved changes to the RCU firmware. The same (or very similar) changes should be tested and implemented for PHOS.

- FEE

There are 28 FEE cards that need to be repaired. Iouri will come to CERN next week to work on that. A RCU is needed for testing the cards.

- AMORE

Amore now works if an environment variable is set correctly. This will be done in the login script once the daq computers are back after the upgrade.

- Shifts

Day and evening PHOS shifts will start from Monday morning.

21 Jan 2010 PHOS Commissioning Meeting

Minutes

- Point 2

A pedestal run was taken on Thursday (14/1) and the pedestal RMSs were identical to what was seen at the end of last year (Mod. 2: 0.93, Mod. 3: 0.84, Mod. 4: 0.70).

Oystein has changed the read-out order from the RCU so that it now alternates between the two branches. This had the effect of reducing the PHOS busy time to below 1 ms and thus enabling running at rates of 1 kHz. This is great progress! It means that we don't have to cut out any PHOS channels to achieve low enough busy time.

- LED system

Yuri and Sasha have worked on the LED system. On module 4, a bad connection to the "black box" on the module was found. After it was reconnected, the LED system again works in Module 4 (except for some well know L-shaped areas). All LEDs in Module 3 can be fired provided that one runs with a single 16x2 band of LEDs at a time. Yuri's presentation is here http://indicobeta.cern.ch/getFile.py/access?contribId=0&resId=1&materialId=slides&confId=81872

- RCU extension on module 3

Bernhard and Sasha have been working on this for the last week, but no solution has been found. It must most likely wait until the module is taken out of the pit.

- FEEs in PVSS

A new Intercom layer is being developed and is currently tested by the TPC and TRD. When it is ready for PHOS, the implementation of control of the PHOS FEEs in PVSS can start. Also the FEE temperatures will then be included in PVSS.

- AMORE

The PHOSagent doesn't work in the latest version. Yuri is working on it.

- Other topics

The weekly SRC meetings will start next Tuesday, 26/2. The daily meetings at Point 2 will start on 1. Feb.

PHOS commissioning in 2009

19 Nov 2009 PHOS Commissioning Meeting

Minutes

- Point 2

The reset of the trigger clock at start of run has been removed. This was the main reason for PHOS going busy earlier. A new ramping sequence for Module 2 has been found. It brings up Module 2 with low noise levels after a power cycle. The ramping sequences are different for Module 2 and Modules 3,4.

- LED

There are problems with the LED system in Module 3. It was decided to reboot the LED system later today at 5 PM.

- APD

Dmitri Peressounko presented some investigations of the noise levels in Module 4 for different voltages and APD settings. See his presentation http://prsnko.web.cern.ch/prsnko/APDsettingsMod4.pdf

- PVSS

Alexander Mamonov listed some of the tasks that should be improved in PVSS: -Add cooling loop status -Update existing interlocks -Update the shift instructions for how to handle interlocks -data to offline database

It was decided, however, to not implement any of these changes at the moment because of the short time expected before the first collisions. The development during next week will probably make the situation clearer and it will then be possible to decide when they can be implemented.

Some settings in PVSS will, however, be changed. The ramping speed when switching off the HV in emergencies will be reduced from 50 V/s to 10 V/s. The maximum allowed temperature for the FEE cards will be reduced from 50 C to 35 C.

12 Nov 2009 PHOS Commissioning Meeting

Minutes

- Point 2

Access yesterday by Iouri and Dmitri. The LED box on Module 3 was replaced. All the TRUs on Module 4 now work. In preparation for the UPS intervention on Wednesday and Thursday the cooling was switched off. It was switched on again after the tests.

- RCU

The noise situation has improved considerably in Modules 2 and 3. The main reason for this is the fix by Iouri and Dmitri of the input LV last week. The pedestal RMS are now 0.92 in Module 2, 0.86 in Module 3, and 1.68 in Module 4. The event sizes in empty events are around 2 kByte/LDC with ZS threshold 3, corresponding to a total PHOS event size of about 12 kByte/event. This is when the latest firmware, which includes zero-skipping, is used on all RCUs.

Øystein found that by turning off one FEE card on the outermost RCU of Module 4 the noise level of that RCU improved considerably. There are still some bad channels in Module 4. Dmitri Peressounko has investigated the behaviour of the bad channels when lowering the HV. The study is still ongoing.

-TRU

The signals from the TRU/TOR should be checked and compared with the LED trigger signals at the CTP.

-ERP tests

PHOS has performed without problems in the last 3 ERP tests. There is still an issue with PHOS going busy. The problem has been solved for the EMCAL, however, by removing the TTCrx reset from the CTP start of run sequence. The same solution should work for PHOS and should be implemented as soon as possible.

5 Nov 2009 PHOS Commissioning Meeting

Minutes

- Point 2

The steel pipes which were dicscovered above Module 2 last week have been rerouted. Several things were done by Iouri and Dmitri during the access yesterday. An attempt was made to reconnect the TRU on the dead branch on Module 4. This was unsuccessful. The LV output was checked and in particular on the 13V channel for Module 3 and 4 large oscillations were detected. The configuration of the Wiener PWS was changed to fix this. The LED box on Module 3 was found to be cold, indicating that it was not working. By disconnecting and reconnecting the cable it came back to life.

- TRU

The right firmware is now installed on all the TRUs on Module 2 and 3. 4 out of the 8 TRUs on Module 2 have been tested with LED signals. Only branch A of the TRUs can be read-out, however. The sending of a PHOS trigger signal to the CTP should be tested and the timing should be checked.

-Pedestals

After the reconfiguration of the Wiener PWS the pedestal RMS is now much lower in Module 3. One does not observe any events where the pedestal baseline is shifted by a large amount from one event to another.

- LED system

Yuri has analysed some LED runs. The output from Module 2 shows some holes with ADC<10 in both the high gain and low gain channels. In Module 3, signals are only seen in less than 1/4 of the module.

29 Oct 2009 PHOS Commissioning Meeting

Minutes

- Point 2

During the access it was discovered that 10 steel pipes with diameter 1 cm had been installed in front of PHOS Module 2. The Technical coordinators have been notified. These pipes have to be rerouted before the start of the run.

The PHOS HV crate turns itself off without warning. It happened around 8 AM yesterday (28/10), at about the same time today (29/10), and at 5 PM today (right after the meeting). It is not clear why this happens. Andre Agustinus will connect the power supply to an UPS to see if that helps.

During the access today, the polarity of the Trigger Level 1 cable with wrong polarity was switched.

- APD

A bug was found by Oystein in the setting of the APDs for module 3. After it was fixed, the noise is lower in Module 3.

- TRU

A bug has been found in the TRU firmware. 3 TRUs in Module 2 now deliver stable triggers in LED runs.

- HLT

The geometrical projection trigger is being tested, but there are memory leaks.

- AOB

Joakim informed about the new PHOS Run Coordinator page https://aliceinfo.cern.ch/Collaboration/Technical_Coordination/ALICE_Project/PHOS/PHOS_shifts/index.html

22 Oct 2009 PHOS Commissioning Meeting

Minutes

- Point 2

As a result of the access yesterday, all RCUs are now installed. The complete Module 3 can be read out, including the bad branch, but the data in that branch appears to be corrupt. The power trips when the TRUs on Module 2 are turned on. One of the L1 cables going from PHOS to CTP has the wrong polarity, which has to be reversed.

- Cooling

The crystals are now at -25 C, the design temperature. When the LV was switched off during the access yesterday, it was noted that the temperature in the warm volume saturated around 5-7 C.

- APD

New APD settings are being calculated by Hisa. They should be for gain 50 at -25 C. The settings for Module 4 should be calculated for 380 V.

- DCS

A fix to leave the HV on when in the state "beam tuning" has been implemented by Kyrre.

- HLT

The geometrical projection trigger is available and ready for testing.

- AOB

Dmitri presented a first list of tasks for shifters.

15 Oct 2009 PHOS Commissioning Meeting

Minutes

- Cooling

Cooling is progressing faster than anticipated last week. The modules are currently at about -7 C and are approaching -15 C. The humidities are 22% in Phos2, 27% in Phos3, 24% in Phos4.

- RCU

One RCU (90) was removed from Module 4 during the access yesterday. It will be tested in the lab and put back next Wednesday.

- HV

As the temperature is now rather rapidly approaching -25 C, the need for APD settings for intermediate temperatures is less urgent. New settings should be calculated for Module 4, however, since it is operating at 380 V (not 395 V).

- DCS

The problems with the intercom layer persist.

- Analysis of data from cosmic and injection test runs

Yuri reported on an analysis of data from the beam injection test 3 weeks ago. The presentation is available from the link below. Boris has looked at the cosmic data and analysed about 3 Mevents. There appears to be some noise also in module 2, see figure below. It was not clear if this amount of noise is what is expected with the observed RMS and threshold settings (For Module 2, pedestal RMS 2, threshold 4).

  • Distribution of clusters in Module 2 for cosmic data (Boris Polichtchouk):
    cluPos.gif

08 Oct 2009 PHOS Commissioning Meeting

Minutes

- Status at Point 2

During the access yesterday (7/10/09) the problem with the LED signals in Module 3 was found and solved. The cable between the racks and the module had to be replaced. The pulse1 LED cable now goes to Module 3.

- Cooling

The reason for the crash last week has been understood and taken care of. A new interlock for the temperature of the compressor has been installed. It is estimated that it will take 2-3 weeks to cool the modules down to -25 C. When the cooling will start has not yet been decided. Next week one hopes to get a better estimate of the current leakage rate.

-RCU

Parts of all modules are now normally included in the global runs during the evening and night shifts. Mod. 2: All RCUs OK. Mod. 3: 3 RCUs OK, but noisy at the moment. Mod. 4: 2 RCUs OK, but one of them has a dead branch.

-HLT

An HLT trigger for tracks entering PHOS is being implemented.

01 Oct 2009 PHOS Commissioning Meeting

Minutes

-Status at Point 2

Last night PHOS ran stable in the global partition with 2 modules (2 and 3, one branch excluded) for 12 h. This is the first time more than one module has been included in the global partition. Module 3 runs fine but is hard to get started after the low voltage has been switched off.

-cooling

Cooling started on Monday (28/9) and +5 C was reached on Tuesday. The humidity in Mod. 4 was then between 25-40%, in Mod. 3 around 22%, and in Mod. 2 12-14%. The temperature remained at +5 C until this morning (1/10) when there was a hardware problem and the cooling was turned off. The problem was fixed and cooling resumed at around 11:00. The modules are now again approaching +5 C.

-RCU

The missing RCU on Mod. 3 was reinstalled during the access on Wednesday (30/9). One of the branches appears OK but the other seems to produce only garbage output. Implementation of zero-skipping on the RCU is under way. This should significantly reduce the PHOS event size.

-HV

Mod. 4 is now running with 380 V. New APD settings should be produced for this voltage. Since we don't know at what temperature the cooling will stop, it would be good to have APD settings for a few different temperatures, for example -5, -10, -15, -20 C.

-DCS

Timur is working on the intercom layer and its inclusion in PVSS. Problems were encountered. An expert on the intercom layer in PVSS should be contacted.

-Trigger

Oystein reported that he can now generate Level-0 triggers in LED runs.

-Analysis of data from beam injection test

Joakim has looked at some of the events from the beam injection test last weekend. One can see what appears to be clear signals when looking at the pulses from individual cells. The amplitudes are always very low, seldom above 10-15 ADC counts in the High Gain channel. The signals also appear later than expected by about 10-15 timebins (1.0-1.5 microsec). Further studies are necessary.

17 Sep 2009 PHOS Commissioning Meeting

Minutes

- cooling

The humidity is going down, but Module 4 (or at least one of its sensors) shows a strange behaviour. Cooling expert should arrive on Sunday.

- RCU

The missing RCU will be installed tomorrow. Bernhard arrived with extension cable. The prospects for implementing sparse read-out on the RCUs are good and test will begin next week.

- HV

The channel that tripped on Module 3 was fixed during the access on Monday. The problem was a broken lemo connection. The HV channels on Module 4 are OK as long as the voltage is below 380 V.

- DCS

Kyrre and Timur are working on debugging the Intercom layer.

- Trigger

A fix has been found for the EmCal trigger which reduces the noise. The same fix should be applicable for PHOS and would improve the trigger noise situation.

- Cosmics

Boris has analysed some 2000 of the events that were recorded in the global physics run with PHOS. No MIP peak has been observed, but that is not expected for that number of events.

- Transfer line test 25-28 September

The current plan is to have PHOS ON during the transfer line test next week. This was done last year without problems. But the potential risks for the detector should be considered before the final decision is made.

03 Sep 2009 PHOS Commissioning Meeting

Minutes

- cooling After the latest magnet access the water cooling is now in a good state. Drying still ongoing.

- RCU

We have 11 DCS boards. One RCU+DCS is now out and is waiting for extention cable. Want to install next Wednesday Zero suppression and sparse readout tests ongoing.

- FEE

As reported before, one branch is dead, and one card is missing on module 4. Everything else seems fine.

- DAQ

Problems with the DDLs on the two lower RCUs on module 4. Investigation together with DAQ.

- DCS

All HW and SW interlocks are in place and displayed in PVSS. Run control unit needs to be fixed to facilitate for different run types (LED system on/off)

- Trigger

One TRU cannot be programmed with new firmware if needed Noisy triggers from TRUs TOR tested together with DAQ TIN server in place.

- HLT

Running with PHOS Online display used for commissioning

- Plan for cosmic data.

When we are ready for running in global partition we will discuss our trigger input. TRU is probably not useful before the modules are cold

- People

Boris will come 1.Sep, Yuri will come to CERN 21.Sep. Hisa will leave 31.Aug and Lijiao will leave on 5.Sep, Misha will leave in the beginning of Octorber. Timur probably come next week.

27 Aug 2009 PHOS Commissioning Meeting

Minutes

- cooling Water cooling leackage was fixed. Now running stably. Humidity is ~10%, ~25%, ~30%.

- RCU We have 11 DCS boards. One RCU+DCS is now out and is waiting for extention cable. One RCU is missing jtag connection, needs test next Monday. New firmware is under testing. New Mapping was commited yesterday.

- FEE As reporeted before, one branch is dead. Another branch is missing connection probably week GTL connection, needs check next Monday. Can't read version from 15 TRUs.

- DAQ DAQ trouble, for example, one missing LDC. We probably can fix with expert's help on surface.

- DCS More hadware&software interlock are visible in PVSS. Different configurations are available. Nice temperature and humidity panels are avalable.

- Trigger TOR's DCS is missing connection and needs to be checked next Monday. Kyrre and Dominik are working integration of the TOR in PVSS.

- HLT Not much progress. Just need data.

- Plan for cosmic data. Some discussions was on e-mail.

- Peoples Boris will come 1.Sep, Yuri will come to CERN 21.Sep. Hisa will leave 31.Aug and Lijiao will leave on 5.Sep, Misha will leave in the beginning of Octorber.

20 Aug 2009 PHOS Commissioning Meeting

Minutes

- Cooling The humidity is decreasing constantly after the dried air (or nitrogen) started five days ago. We probably need to wait for 3 weeks in order to fix water cooling leackage because of L3 magnet.

- DCS Svain and Alexander has worked since last week and PVSS is close to what we need now. DCS for LED system is running successfully. New cooling panel in PVSS. We need the state machine for FEE systems, which Timor can handle in the future.

- RCU RCU firmware v3 is now working. There happens a corraption of ALTRO data for both TPC and us, which experts are working for. 8 DCS boards needs to be reinitialized.

- FEE New mapping are on trunk.

- DAQ 1.2kHz with busy box is promissing.

- HLT Two triggers, high energy threshold trigger and mip trigger, are implemented, tested, and submitted by Svain.

- Trigger TRU+TOR is going to L0. L1 needs to be tested in the future by Lijiao and Yusuke.

- Plan for cosmic runs Let's wait for cooling. Need input how many days of cosmic runs is needed for calibration.

- Man-power at CERN Timor is at Oslo now and will come to CERN soon but not next week. Bernhard, Oystein, Svain, Lijiao, Sasha, Misha, Alexander, Yaping, Fumihiro, Yusuke, Hisa are at CERN now. Svain will leave at the end of next week. Fumihiro will leave this Saturday. Hisa will leave at the end of August. Lijiao will leave on 5th/September. Yuri and Boris has called in.

13 Aug 2009 PHOS Commissioning Meeting

Agenda

Minutes

- DCS Svain and Alexander need to work together. Oystein will come to CERN and arrange the job.

- FEE Mapping for DCS is not yet fixed. New RCU firmware work. The sparse readout must be tested soon. DAQ integration test is planned tomorrow, this Friday. Trigger things need to be tested.

- DAQ Need test high trigger DAQ next week.

- Cooling Cooling pipes were disconnected for installation of PHOS#2&1

- Plan for cosmic runs. We are ready. Need to discuss about shifts

- HLT Creat a componenet of energy threshold trigger and will ask to implement the working one.

- PHOS#1&2 PHOS#1 was transferred to P2 and will be installed tomorrow. PHOS#2 is being installed now!!! Let's hear more update from Sasha. PHOS#1 at build167 had 2 bad FEEs and 1 bad branch.

- Peoples Oystein will reach at the Geneve this afternnon. Svain will be at CERN for 2-3 weeks depending on the update. Yuusuke has just come to CERN and stay for 2 months. Hiro will leave CERN on 22.08 and Hisa will leave on 31.08. Yaping is at CERN.

6 Aug 2009 PHOS Commissioning Meeting

Agenda

Minutes

- DCS
Kyrre and Alexander need to communicate remotely what's done before.

- FEE
Mapping between DDL to DAQ are good, but the mapping for DCS is not yet fixed and will be fixed this afternoon by Svein. New RCU firmware are not working for both PHOS and EMCAL. Need update about the PHOS trigger. DAQ integration is planed on Friday next week, and then trigger integration test will follow. PHOS#1 has 32 bad channels including 2 suspicious T-cards and all TRU are working

- DAQ
Need test >200Hz with busy box. We could read 1.4kHz before.

- Cooling
Cannot start cooling down before at least mid-September.

- Plan for cosmic runs
Basically we are ready for cosmic runs. Need to discuss about shifts.

- HLT
Need proper mapping.

- PHOS#2 schedule
Will be installed after extra 5 days.

- Peoples Schedule
Timor will start for the database in three weeks.
Oystein will be absent from tomorrow for a week.
Dong and Lijiao has left CERN.
Kyrre will come to CERN 10.08 and Alexander will leave CERN on 13.08, so both have overlapp.
Yaping will continue the trigger works.
Hiro will leave CERN on 22.08 after PHOS#1 installation.

29 July 2009 PHOS Commissioning Meeting

Agenda

I propose that we follow Yuri's list for the discussion (with some additions):

1) DCS: a list of DCS integration tasks was presented in Serguei's talk at the last PHOS meeting, Kyrre was working on it in July, today the DCS experts from Sarov are expected to arrive to CERN. Tomorrow the DCS integration session should be held with PHOS in ACR. How far is PHOS from the full integration?

2) FEE: what is the status of the FEE configuration in detector configuration DB, and the procedures to load the configuration from it? The RCU firmware v.3 should be tested and compared with similar data taken with the old RCU firmware. RCU trailer registers could be discussed and loaded to RCU. Shall we run with zero suppression by default? The status of the mapping of the channels? Bad channel maps?

3) TRU: is it ready for MIP selection? Can it be already used for cosmic run with PHOS as a trigger detector? Does Dong et al. need any special LED runs to set TRU thresholds and active channel list?

4) DAQ: online monitoring in AMORE can be tested again, as soon as DAQ experts fix the problem with PHOS QA module.

5) ECS: are we happy with 2 standalone run types (PEDESTAL and LED), or any additional run types are needed? For example, does TRU need any dedicated run types? Besides, ECS should be equipped by a script to stop LED after the LED run, otherwise LED will continue flashing for the next runs.

6) Colling: when PHOS can be cooled down to -25 degrees? Will it be ready before the cosmic run which starts on 22 August?

7) Plans for cosmic runs: we definitely should participate in it, either worm or cold. One has to foresee the breaks between the physics runs to have LED runs (at least, once per day, but better every shift). We could discuss what is needed for the full-dress rehearsal of the data processing, starting from the start of run and finishing the raw data reconstruction and OCDB filling with the calibration parameters.

Minutes

See Indico page

As we agree, we continue the (definitely short) PHOS commissioning meeting every Thursday from 10am. Here is the minutes for today, 29.07.09

- DCS The DCS integration test finished succesfully yesterday, on 28.07. Some improvement is necessary, but our system passed their criteria. Some issues; Specific LV-power-up procedure is necessary in PVSS. Central message connection to ECS system is needed. Presentation at DCS meeting this afternoon.

Kyrre is working at CERN. Alexander Mamonov has just arrived at CERN.

- FEE&RCU The configuration data base basically exists Some information must be transferred in OCDB, which is not in highest priority. RCU firmware v3 doesn't work, doesn't take any data from FEE properly. An expert will come to help us. Zero suppression mode is default for physics runs. During commissioning and cosmic runs, we should take with/without zero suppression, someone should compare them. Some system (DA and online monitoring) needs zero suppression flag, zero suppression offset and zero suppression threshold, Oystein tested zero suppression info into RCU trailer. What we have in current RCU trailer; zero suppression flag, total number of samples, number of readout samples, sampling frequency. Dieter's comment; the zero suppression offset and threshold must be fixed through all runs once we know the noise situation. New channels mapping is coming for FEE, and is ready for TRU.

Timor is supposed to come to CERN for PHOS for finish the configuration database.

- TRU One PHOS trigger is already in the global trigger system. We need test with LED flushing. The 2x4 LED channels is the minimum LED unit. TRU and TOR needs the feeserver upgrade.

Dong has left CERN and Lijiao will be leaving today. Need man-power. Timor plans to work for the feeserver.

- DAQ We couldn't take >200Hz even with zero-suppression. DAQ provided 6 LDCs instead of 3 LDCs, we will see the limitation tomorrow. Once ECR people started AMORE, someone should test the online monitoring again. Need DAQ integration test after DCS PVSS implementation.

Need man-power.

- HLT Clusterization and inputting into ESD is working 1kHz. A PHD student, Svein Lindal, started to implement the trigger feature. Online monitoring of QA is not yet connected to AMORE.

- ECS Additional run modes in addition to LED and PEDESTAL runs is not needed. PVSS should control on/off the LED system, so such a physics run with LED flushing should not occur.

- Cooling In phos meeting this afternoon.

- Plan for cosmic runs PHOS is not declared as a part of global partition during global cosmic runs. We have to be part of global runs. What we need at first is to test in standalone mode.

Let's discuss in phos meeting this afternoon and in e-mails. Oystein will contact Jan.

- PHOS#2&1 at PHOS lab Noise is surprisingly small. PHOS#2 is basically ready for installation.

Yaping and Fumihiro will take electrical test for PHOS#2.

PHOS commissioning in 2007

27 Nov 2007 / J. Alme / Electronics / DCS and Actel firmware upgrade

The upgrade of the DCS board firmware to v2.8 and Actel fw to v1.4 went without problems yesterday on RCUs of the 1st PHOS module. The reason for the upgrade is bug that makes the RCU Xilinx be reprogrammed from the RCU Flash whenever the DCS board reboots. More details concerning the new versions can be found at the Bergen ALICE Wiki.
Note that a script (S31_setOldMode) in /etc/boot/ has been added to make the DCS firmware backwards compatible with the current RCU Xilinx firmware. This script must be removed when the new RCU firmware is ready.

22 Nov 2007 / B. Skaali / Electronics / GTL strips

Tests on a revised GTL strip will start in beginning of December in Oslo. The exact date can not be given today due to high work load together with illness at the Electronics Lab.

Welcome to the PHOS commissioning wiki page

-- Created by CesarMatos - 22 Nov 2007
Topic attachments
I Attachment History Action Size Date Who Comment
PowerPointppt PHOSinPhysicsRuns.ppt r1 manage 500.5 K 2009-10-15 - 17:43 JoakimNystrand Analysis of data from beam injection test (Yuri Kharlov)
GIFgif cluPos.gif r1 manage 16.0 K 2009-10-15 - 17:46 JoakimNystrand Distribution of clusters in Module 2 for cosmic data (Boris Polichtchouk)
Edit | Attach | Watch | Print version | History: r35 < r34 < r33 < r32 < r31 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r35 - 2010-11-13 - YuriKharlov
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    AlicePhos All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright &© 2008-2023 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback